OpenStudy (loser66):

Help calculate the area of the graph of f(x,y) = 3/2

OpenStudy (loser66):



OpenStudy (loser66):

The total area = area ABED + area BCE area ABED = \((\dfrac{1}{\sqrt2}-\dfrac{1}{2})(\dfrac{1}{2})=0.10355\)

OpenStudy (anonymous):

You can determine the area with a double integral: \[\int_{1/2}^1\int_{1/2}^{\sqrt x}dy~dx\]

OpenStudy (loser66):

yes, but I can't find out the right answer from it

OpenStudy (loser66):

After calculating, I got 0.116 for this part, adding with the area of rectangle, I got 0.219 while the answer from my prof is 0.2714 he said it is \(\dfrac{5}{8}-\dfrac{1}{2\sqrt 2}=0.2714\) and asked us find out the way to get that solution :(

OpenStudy (loser66):

oh, no, it is not your int

OpenStudy (ageta):

.2714 is wrong

OpenStudy (loser66):

B =(\sqrt 2/2, 1/2)

OpenStudy (ageta):

yes

OpenStudy (ageta):

\[\int\limits_{1}^{9}\]

OpenStudy (loser66):

hence, the int is \[\int_{1/2}^1\int_{1/\sqrt 2}^{\sqrt y}3/2dx~dy\]

OpenStudy (loser66):

I check on wolfram and it has the same answer with me, how to get 0.2714??

ganeshie8 (ganeshie8):

you could also find the area and multiply by 3/2 as the function is a constant

OpenStudy (loser66):

Question: why do we have to take the first integral when it is a concrete rectangular?

ganeshie8 (ganeshie8):

You don't need to use double integrals if you don't want to as the function is constant

ganeshie8 (ganeshie8):

\[\iint_R f(x, y) dA = \iint_R 3/2 dA = 3/2 (\text{Area})\]

ganeshie8 (ganeshie8):

You're integrating the function f(x,y) = 3/2 over the given region. This is a volume problem :